# Heterogeneous Treatment Effects and

### v3

@@ Appendix A Additional Classification of Articles
Table A.1 provides an additional classification of the articles described in Table 1, by research design. Note that we collapse difference-in-difference and panel analyses into one category that includes two-way fixed-effects and other estimators of the average treatment effect on the treated (ATT). We also collapse IV and natural experimental analyses into a single category that includes studies with some claim of exogenous variation not created by researchers that is argued to facilitate identification of an average treatment effect (ATE); an intent-to-treat effect (ITT); or a local average treatment effect (LATE). [^18] This LATE is often termed the complier average causal effect (CACE) in political science. This table shows that using HTEs to detect mechanisms is not unique to any one research design in common usage; the proportions of articles in these journals that uses HTEs as a mechanism test (given by the “weighted average” column) is quite similar across all of these designs.
@@
Appendix B Motivating Example
B.1 Incorrect DAG
Figure A.1 depicts the DAG that is evaluated by the HTE analysis in Remarks 1-2. Note that the dashed lines are absent in the theoretical model. In the model, only the learning mechanism is active. This mechanism is evaluated by examining heterogeneity in treatment effects with respect to voters’ prior beliefs.
We note that this graph does not precisely correspond to the model in the paper. Here, the valence shock, $v_{i}$ is measured pre-treatment (ex-ante), and the researcher (wrongly) believes that it moderates treatment effects. We represent this in the graph with $\widetilde{v}_{i}$, a measure of “ex-post” valence.
@@ B.1 Incorrect DAG
Figure A.1: Incorrect directed acyclic graph representation (relative to the model). The dashed arrows are not implied by the model but correspond to a test of the valence mechanism.
B.2 Proofs of Remarks 1-2
@@
$CATE(y_{1},X=\pi)$ $=E[y_{1}|\omega^{\prime\prime},\pi]-E[y_{1}|\omega^{\prime},\pi]$ (16)
$=E[\beta(\overline{\theta}|\pi,\omega^{\prime\prime})+v_{i}|\omega^{\prime\prime},\pi]-E[\beta(\overline{\theta}|\pi,\omega^{\prime})+v_{i}|\omega^{\prime},\pi]$ (17)
$=E[\beta(\overline{\theta}|\pi,\omega^{\prime\prime})]-E[\beta(\overline{\theta}|\pi,\omega^{\prime})]+E[v_{i}]-E[v_{i}]$ (18)
$=E[\beta(\overline{\theta}|\pi,\omega^{\prime\prime})]-E[\beta(\overline{\theta}|\pi,\omega^{\prime})]$ (19)
where equation (17) to (18) follows from the linearity of expectations. Similarly, we have:
$CATE(y_{1},X=\pi^{\prime})=E\beta(\overline{\theta}|\pi^{\prime},\omega^{\prime\prime})-E\beta(\overline{\theta}|\pi^{\prime},\omega^{\prime})$
Recall that $\beta(\overline{\theta}\mid\pi_{i},\omega)$ is given by:
$\beta(\overline{\theta}|\pi_{i},\omega)$ $=\frac{\pi_{i}\phi(g-f(\overline{\theta},\omega))}{\pi_{i}\phi(g-f(\overline{\theta},\omega))+(1-\pi_{i})\phi(g-f(\underline{\theta},\omega))}=\frac{1}{1+\frac{1-\pi_{i}}{\pi_{i}}\frac{\phi(g-f(\underline{\theta},\omega))}{\phi(g-f(\overline{\theta},\omega))}}$
Given function $f$ and pdf $\phi$, we conclude $CATE(y_{1},X=\pi)\neq CATE(y_{1},X=\pi^{\prime}).$
@@
$CATE(y_{1},v)$ $=E[y_{1}|\omega^{\prime\prime},v]-E[y_{1}|\omega^{\prime},v]$ (20)
$=E[\beta(\overline{\theta}|\pi,\omega)+v|\omega^{\prime\prime},v]-E[\beta(\overline{\theta}|\pi,\omega)+v|\omega^{\prime},v]$ (21)
$=E[\beta(\overline{\theta}|\pi,\omega^{\prime\prime})]-E[\beta(\overline{\theta}|\pi,\omega^{\prime})]$ (22)
$=E[\beta(\overline{\theta}|\pi,\omega^{\prime\prime})+v^{\prime}|\omega^{\prime\prime},v^{\prime}]-E[\beta(\overline{\theta}|\pi,\omega^{\prime})+v^{\prime}|\omega^{\prime},v^{\prime}]$ (23)
$=E[y_{1}|\omega^{\prime\prime},v^{\prime}]-E[y_{1}|\omega^{\prime},v^{\prime}]$ (24)
$=CATE(y_{1},v^{\prime})$ (25)
Note that $v$ is independent of $\omega$. As a result, equality holds from (22) to (23) when we add $v^{\prime}$ to both expectations. ∎
@@
$y_{2}$ =$=\begin{cases}1&\text{ if }\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}-\pi^{C}\geq 0\\ 0&\text{ else }\end{cases}$ (26)
$CATE(y_{2},X=\pi)$ is therefore given by:
$CATE(y_{2},m=\pi)$ $=E[y_{2}|\pi,\omega^{\prime\prime}]-E[y_{2}|\pi,\omega^{\prime}]$
$=\Pr[y_{2}=1|\pi,\omega^{\prime\prime}]-\Pr[y_{2}=1|\pi,\omega^{\prime}]$
$=\Pr[\beta(\overline{\theta}|\omega^{\prime\prime},\pi)+v_{i}-\pi^{C}\geq 0]-\Pr[\beta(\overline{\theta}|\omega^{\prime},\pi)+v_{i}-\pi^{C}\geq 0]$
$=\Pr[v_{i}\geq\gamma(\pi,\omega^{\prime\prime})]-\Pr[v_{i}\geq\gamma(\pi,\omega^{\prime})],$
where $\gamma(\pi,\omega)=\pi^{c}-\beta(\overline{\theta}|\omega,\pi).$ Note that $\gamma(\pi,\omega)\in[-1,1]$. Because $v_{i}\sim U(-1,1)$, and the posterior $\beta(\overline{\theta}|\omega,\pi)$ is continuous in $\pi$, $CATE(y_{2},\pi)-CATE(y_{2},\pi^{\prime})\neq 0$ almost everywhere. ∎
@@
$CATE(y_{2},v)$ $=E[y_{2}|v,\omega^{\prime\prime}]-E[y_{2}|v,\omega^{\prime}]$
$=\Pr[y_{2}=1|v,\omega^{\prime\prime}]-\Pr[y_{2}=1|v,\omega^{\prime}]$
$=\Pr[\beta(\overline{\theta}|\omega^{\prime\prime},\pi_{i})+v-\pi^{C}\geq 0]-\Pr[\beta(\overline{\theta}|\omega^{\prime},\pi_{i})+v-\pi^{C}\geq 0]$
To calculate the above probability, the randomness comes from $\pi_{i}$. It is useful to rewrite $\beta(\overline{\theta}|\omega,\pi_{i})+v_{i}-\pi^{c}\geq 0$ so that we can separate $\pi_{i}$ and other non-random components:
$\frac{\pi_{i}}{1-\pi_{i}}\geq\frac{\phi(g-f(\underline{\theta},\omega))}{\phi(g-f(\overline{\theta},\omega))}\frac{\pi^{c}-v}{1-\pi^{c}+v}$ (27)
Note that $\frac{\pi_{i}}{1-\pi_{i}}$ is monotone in $\pi_{i}$, which has distribution $F_{\pi}$. We use $\alpha(\omega,v_{i})$ to denote the RHS of (27). We can then express $\Pr(\beta(\overline{\theta}|\omega,\pi_{i})+v-\pi^{C})$ as $F_{\pi}[\frac{\alpha(\omega,v)}{1+\alpha(\omega,v)}]$, so the CATE is given by:
$CATE(y_{2},v)=F_{\pi}[\frac{\alpha(\omega^{\prime\prime},v)}{1+\alpha(\omega^{\prime\prime},v)}]-F_{\pi}[\frac{\alpha(\omega^{\prime},v)}{1+\alpha(\omega^{\prime},v)}]$
It is clear that $CATE(y_{2},v)$ depends on the values of $v$ and $\alpha$. If there exists at least one $\alpha>0$ so that $\frac{\alpha}{1+\alpha}\in(0,1)$, then we can easily find $CATE(y_{2},v)\neq CATE(y_{2},v^{\prime})$. A sufficient condition for $\alpha\in(0,1)$ is $\min\{v,v^{\prime}\}<\pi^{C}$.
@@
Appendix C Generalized Framework with Multiple Mechanisms
@@
$TE(z,z^{\prime};X)$ $=Y(z,M_{1}(z;X),...,M_{J}(z;X);X)-Y(z^{\prime},M_{1}(z^{\prime};X),...,M_{J}(z^{\prime};X);X)$ (28)
@@
$DE(z,z^{\prime};X)=Y(z,M_{1}(z;X),...,M_{J}(z;X);X)-Y(z^{\prime},M_{1}(z;X),...,M_{J}(z;X);X)$ (29)
$IE_{j}(z,z^{\prime};X)=Y$ $(z^{\prime},M_{j-}(z^{\prime};X),M_{j}(z;X),M_{j+}(z;X);X)-$ (30)
@@
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediators at potential outcomes $M_{j}(z;X)$. It is not necessary to believe that treatments produce unmediated (direct) effects on outcomes to use this framework. We allow for direct effects in the interest of generality. The indirect effects, measures the effect on the outcome that operates by changing the potential outcome of the mediator. As is standard, we can then re-write the total effect as follows:
$TE(z,z^{\prime};X)$ $=DE(z,z^{\prime};X)+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X),$ (31)
@@
$ATE(z,z^{\prime})$ $=E_{X}[Y(z)-Y(z^{\prime})]$ (32)
$=E_{X}[DE(z,z^{\prime};X)+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X)]$ (33)
$=ADE(z,z^{\prime})+\sum_{j=1}^{J}AIE_{j}(z,z^{\prime})$ (34)
We use $ADE$ and $AIE$ to denote average direct effect and average indirect effect.
Because there are multiple mediators, we require a further assumption beyond 2 (reproduced from the main text below). Assumption 3 rules out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$.
Assumption 2 (Exclusion I).
Given $z,z^{\prime}\in Z$ and $x,x^{\prime}\in X_{k}$, $X_{k}$ is excluded to the direct effect such that $ADE(z,z^{\prime};X_{k}=x)=ADE(z,z^{\prime};X_{k}=x^{\prime})$.
@@
Given $z,z^{\prime}\in Z$ and $x,x^{\prime}\in X_{k}$, $X_{k}$ is excluded to the indirect effect of any other mechanism, $j^{\prime}\neq j$, $IE_{j^{\prime}}$, if: $AIE_{j^{\prime}}(X_{k}=x)=AIE_{j^{\prime}}(X_{k}=x^{\prime})$.
Assumptions 2-3 constrain the relationship between a moderator, $X_{k}$, and other mechanisms/the direct effect. Figure A.2 depicts violations of Assumptions 2-3 graphically for a MDV ($X^{MDV}$) of mechanism 1, $M_{1}$. Assumption 2 rules out the blue dashed path from $X^{MDV}$ to the direct effect of $Z$ on $Y$. Assumption 3 rules out either/both red dot-dashed path from $X^{MDV}$ to the effect of $Z$ on a different mechanism, $M_{2}$. This means that $X_{k}$ does not moderate the effect of $Z$ on $M_{2}$. Nor does $X_{k}$ moderate the effect of $M_{2}$ on $Y$. However, there can be a direct arrow from $X_{k}$ to $M_{2}$. This means that $X_{k}$ can cause (or predict) the level of $M_{2}$, but it cannot interact with treatment in any way to cause $M_{2}$ or the effect of $M_{2}$ on $Y$. These assumptions are crucial for the use of HTEs to detect mechanisms. In their absence, we cannot link a heterogeneous treatment effect to a unique theoretical mechanism representation (mediator) because we cannot ensure additive separability of the effect of the mechanism of interest from the effects of other (possible) mechanisms.
[…]
Figure A.2: Assumption 2 rules out the blue dashed path. Assumption 3 rules out both of the red dot-dashed paths. All orange and black solid paths are permissible under Assumptions 2 and 3. For $X_{k}$ to be an MDV, at least one of the two orange paths must be present.
Appendix D Directly/Indirectly-Affected versus Latent/Observed Outcomes
We propose a distinction between directly and indirectly-affected outcomes. This distinction is a theoretical claim about the causal process that generates outcomes of interest. This is different from the distinction between latent and observed outcomes. The distinction between latent and observed outcomes invokes both a theoretical claim about the causal process linking latent to observed outcomes as well as an empirical claim about the (non)-observability of the latent outcome. To show that these are distinct classifications, we propose three hypothetical studies united by a common mechanism—learning—in different contexts and organize outcomes into Table A.2. Note that we index outcomes by their sequence in the subscript and across the hypothetical studies $S1,S2,$ or $S3$ in the superscript.
  • Study 1: [Inspired by Coppock (2022) and related information experiments.] A researcher randomly assigns a subject to receive information on a policy issue (e.g., gun control) or a placebo message. They then measure a subject’s attitudes about gun control policy proposals on a Likert scale. The proposed mechanism (learning) affects attitudes, which are assumed to be latent, which in turn shape responses to a (discrete) Likert scale.
    • $y_{1}^{S1}$: Attitudes about gun control policy
    • $y_{2}^{S1}$: Likert-scale approval for gun control policy
  • Study 2: [Inspired by motivating example.] Nature stochastically assigns subjects to an adverse shock (e.g., a natural disaster). Researchers conduct a survey after-the-fact measuring both willingness to pay for the incumbent (relative to a challenger) and intention to vote for the incumbent. The proposed mechanism (learning) affects beliefs about the incumbent’s type, which affects expected utility (measured by willingness-to-pay) and then vote choice (a non-linear function of utility). The outcomes are thus:
    • $y_{1}^{S2}$: Willingness to pay for incumbent
    • $y_{2}^{S2}$: Intended vote choice for incumbent
  • Study 3: Suppose researchers conduct an experiment in which they provide information about a proposed policy to legislators in an assembly that could use either recorded and voice votes to pass a bill. The informational intervention affects researcher beliefs about the ideal point of the proposed policy (which they compare to a known status-quo) through a learning mechanism. These beliefs shape assesments of expected utility of the bill relative to the status-quo, which in turn shape vote choice. However, the observability of an individual legislator’s vote depends on the voting method used to pass the bill (recorded or voice vote), as follows:
    • $y_{1}^{S3}$: Expected utility of bill (relative to status-quo)
    • $y_{2A}^{S3}$: Voice vote on bill (individual votes are not observed)
    • $y_{2B}^{S3}$: Recorded vote (individual votes are observed)
Table A.2 organizes the outcomes for these hypothetical studies according to their classification as directly/indirectly-affected and latent/observable outcomes. As there are outcomes that fall in all four cells, these examples illustrate that these concepts are distinct.
Latent (Unobserved/Unobservable) Observed
[…]
Table A.2: Classification of outcomes for three hypothetical studies indicates that the directly-/indirectly-affected classification of outcomes is distinct from the distinction between latent and observable variables.

### v4

@@ Appendix A Additional Classification of Articles
Table A.1 provides an additional classification of the articles described in Table 1 but disaggregates by research design. Note that we collapse difference-in-difference and panel analyses into one category that includes two-way fixed-effects and other estimators of the average treatment effect on the treated (ATT). We also collapse IV and natural experimental analyses into a single category that includes studies with some claim of exogenous variation not created by researchers that is argued to facilitate identification of an average treatment effect (ATE); an intent-to-treat effect (ITT); or a local average treatment effect (LATE). [^20] This LATE is often termed the complier average causal effect (CACE) in political science. This table shows that using HTEs to detect mechanisms is not unique to any one research design in common usage; the proportions of articles in these journals that uses HTEs as a mechanism test (given by the “weighted average” column) is quite similar across all of these designs.
@@
Appendix B Examples of Current Practice
Appendix C Motivating Example
C.1 Data Generating Process
Figure A.1 depicts the data generating process that is evaluated by the HTE analysis in Remarks 1-2. In the model, only the distaste for corruption mechanism is active. This mechanism is evaluated by examining heterogeneity in treatment effects with respect to $\lambda$. It is clear that the partisan alignment does not affect the distaste for corruption mechanism.
@@ C.1 Data Generating Process
Figure A.1: Data generating process of the model of corruption revelation and behavior. The arrow from the $\lambda$ to the path between the information and the mediator indicates that the $\lambda$ moderates the distaste for corruption mechanism.
C.2 Proofs of Remarks 1-2
@@
$CATE(y_{1},X_{1}=\lambda)$ $=E[y_{i1}|c_{i}=1,\lambda]-E[y_{i1}|c_{i}=0,\lambda]$
$=E[-\lambda+a_{i}-\varepsilon_{i}]-E[a_{i}-\varepsilon_{i}]$
$=-\lambda$
and
$CATE(y_{1},X_{1}=\lambda^{\prime})$ $=E[y_{i1}|c_{i}=1,\lambda]-E[y_{i1}|c_{i}=0,\lambda]$
$=E[-\lambda^{\prime}+a_{i}-\varepsilon_{i}]-E[a_{i}-\varepsilon_{i}]$
$=-\lambda^{\prime}$
Because $\lambda>\lambda^{\prime}$, we conclude $|CATE(y_{1},X_{1}=\lambda)|>|CATE(y_{1},X_{1}=\lambda^{\prime})|.$
@@
$CATE(y_{1},a)$ $=E[y_{i1}|c_{i}=1,a]-E[y_{i1}|c_{i}=0,a]$
$=E[-\lambda_{i}+a-\varepsilon_{i}]-E[a-\varepsilon_{i}]$
$=-\mathbb{E}[\lambda_{i}]$
$CATE(y_{1},a^{\prime})$ $=E[y_{i1}|c_{i}=1,a^{\prime}]-E[y_{i1}|c_{i}=0,a^{\prime}]$
$=E[-\lambda_{i}+a^{\prime}-\varepsilon_{i}]-E[a^{\prime}-\varepsilon_{i}]$
$=-\mathbb{E}[\lambda_{i}]$
As a result, we conclude $CATE(y_{1},X_{2}=a)=CATE(y_{1},X_{2}=a^{\prime}).$
@@
[…]
$CATE(y_{2},X_{1}=\lambda)$ is therefore given by:
$CATE(y_{2},X_{1}=\lambda)$ $=E_{a_{i},\varepsilon_{i}}[y_{i2}|\lambda,c_{i}=1]-E_{a_{i},\varepsilon_{i}}[y_{i2}|\lambda,c_{i}=0]$
$=\mathbb{E}_{a_{i}}[\Pr(\varepsilon_{i}\geq\lambda-a_{i})]-\mathbb{E}_{a_{i}}[\Pr(\varepsilon_{i}\geq-a_{i})]$
$=\sum_{a_{i}\in\{-1,0,1\}}\frac{1}{3}[\Phi(-a_{i})-\Phi(\lambda-a_{i})]$
the second line follows the independence of $\varepsilon$, and $\Phi$ is the standard normal CDF.
Similarly,
$CATE(y_{2},X_{1}=\lambda^{\prime})$ $=\sum_{a_{i}\in\{-1,0,1\}}\frac{1}{3}[\Phi(-a_{i})-\Phi(\lambda^{\prime}-a_{i})]$
Therefore, $CATE(y_{2},X_{1}=\lambda)-CATE(y_{2},X_{1}=\lambda^{\prime})=\frac{1}{3}\sum_{a_{i}\in\{-1,0,1\}}[\Phi(\lambda^{\prime}-a_{i})-\Phi(\lambda-a_{i})]<0$ when $\lambda>\lambda^{\prime}$. Recall the treatment effect is negative (for example from the proof of Remark 1(b)). Therefore the result holds.
@@
$CATE(y_{2},a)$ $=E_{\lambda}[y_{i2}|a,c_{i}=1]-E_{\lambda}[y_{i2}|a,c_{i}=0]$
$=\Phi(-a)-E_{\lambda}[\Phi(\lambda_{i}-a)]$
Now, we compare $CATE(y_{2},0)$ and $CATE(y_{2},1)$. Firstly, we define $g(\lambda)=\Phi(\lambda)-\Phi(\lambda-1)+\Phi(-1)-\Phi(0)$.
Note that when $0<\lambda<1$, $g(\lambda)>0$. This can be seen from $g(0)=g(1)=0$, and it increases ($g^{\prime}(\lambda)>0$) when $0<\lambda\leq 0.5$, and decrease $($$g^{\prime}(\lambda)<0$)) when $0.5\leq\lambda<1$.
Therefore,
$CATE(y_{2},1)-CATE(y_{2},0)=\mathbb{E}_{\lambda}[g(\lambda)]>0.$
Recall the treatment effect is negative (for example from the proof of Remark 1(b)). As a result, $|CATE(y_{2},1)|<|CATE(y_{2},0)|$.
Next, we compare $CATE(y_{2},1)$ and $CATE(y_{2},-1)$. Define $h(\lambda)=\Phi(\lambda+1)-\Phi(\lambda-1)+\Phi(-1)-\Phi(1)$. Note that $\frac{d}{d\lambda}[\Phi(\lambda+1)-\Phi(\lambda-1)]=\phi(\lambda+1)-\phi(\lambda-1)<0$ for $\lambda>0$. Therefore, $h(\lambda)<0$ for $\lambda>0$. Then,
$CATE(y_{2},1)-CATE(y_{2},-1)=\mathbb{E}_{\lambda}[h(\lambda)]<0.$
[…]
@@
C.3 Simulation
We conduct a simulation based on data generating processes based on our motivating theoretical model. Specifically, we generate datasets with $n\in\{100,500,1000\}$ observations under the following parametric and distributional assumptions:
  • $c_{i}\in\{0,1\}$ and $\Pr(c_{i}=1)=\frac{1}{2}$
  • $a_{i}\in\{-1,0,1\}$ where $\Pr(a_{i}=-1)=\Pr(a_{i}=0)=\Pr(a_{i}=1)=\frac{1}{3}$
  • $\lambda\sim U[0,1]$
  • $\varepsilon\sim\mathcal{N}(\mu,1)$ for $\mu=\Phi^{-1}(0.1,0.2,...0.9)+0.25$, where $\Phi^{-1}(\cdot)$ is the inverse CDF of the standard normal distribution. We add $E[c_{i}\lambda_{i}+a_{i}]=0.25$ to generate simulated electorates with 10%, 20%,…,90% incumbent support (in expectation). By varying $\mu$, therefore, we vary the ex-post vote-share for the incumbent that the researcher would observe after running the experiment.
We generate outcome variables as in the main text:
$y_{i1}$ $=-c_{i}\lambda_{i}+a_{i}+\varepsilon_{i}$
$y_{i2}$ $=I[y_{i1}\geq 0]$
Given each simulated data set, we estimate the following OLS model:
$y_{il}$ $=\beta_{0}+\beta_{1}c_{i}+\beta_{2}\lambda_{i}+\beta_{3}a_{i}+\beta_{4}c_{i}\lambda_{i}+\beta_{5}c_{i}a_{i}+\epsilon_{i},$ (14)
for $l\in\{1,2\}$ (both outcomes). We are interested in estimates of $\beta_{4}$ and $\beta_{5}$. $\beta_{4}$ is the estimator of the difference in CATEs moving from the minimum to the maximum of corruption aversion, $\lambda$. $\beta_{5}$ is the estimator of the difference in CATEs for a one category shift in partisanship (e.g., opposition-aligned to non-aligned or non-aligned to aligned).
Figure A.2 reports the average estimates of the differences in CATEs ($\widehat{\beta_{4}}$ and $\widehat{\beta_{5}}$ from (14)). Figure A.3 reports the power of the estimators $\beta_{4}$ and $\beta_{5}$ at the $\alpha=0.05$ level. Since inferences about mechanisms are typically made on the basis of whether we detect HTEs (not the magnitude of the difference), Figure A.3 corresponds to more typical use of HTEs. Note that for all sample sizes, in highly imbalanced electorates (i.e., average incumbent support of 10% or 90%), the $\beta_{5}$ is a better powered estimator than $\beta_{4}$.
Figure A.2: Difference in CATEs in each covariate ($\lambda_{i}$ and $a_{i}$) for outcomes $y_{i1}$ and $y_{i2}$. Each point reports the average over 5,000 simulated datasets.
Figure A.3: Power of the difference in CATEs in each covariate ($\lambda_{i}$ and $a_{i}$) for outcomes $y_{i1}$ and $y_{i2}$. Each point reports the average over 5,000 simulated datasets.
Given the discrete distribution of $a_{i}$, researchers may opt for a more flexible regression specification than (14). To this end, we also consider the following model that compares effects of corruption revelation among independent/neutral ($a_{i}=0$) and aligned ($a_{i}=1$) partisans relative to its effects among unaligned ($a_{i}=-1$) partisans:
$y_{il}=\beta_{0}$ $+\beta_{1}c_{i}+\beta_{2}\lambda_{i}+\beta_{3}\mathbb{I}(a_{i}=0)+\beta_{4}\mathbb{I}(a_{i}=1)+\beta_{5}c_{i}\lambda_{i}+\beta_{6}c_{i}\mathbb{I}(a_{i}=0)+\beta_{7}c_{i}\mathbb{I}(a_{i}=1)+\epsilon_{i}.$ (15)
Figure A.4 plots average estimates of the differences in CATEs ($\widehat{\beta_{5}}$, $\widehat{\beta_{6}}$, and $\widehat{\beta_{7}}$ in (15)) while Figure A.5 plots the power of these estimators at the $\alpha=0.05$. We see that particularly in highly imbalanced electorates, researchers are as likely to detect differences in the CATEs between unaligned (base category, $a_{i}=-1$) and aligned ($a_{i}=1$) voters as they are to detect differences in CATEs in corruption aversion ($\lambda_{i}$).
Figure A.4: Difference in CATEs in each covariate ($\lambda_{i}$ and $a_{i}$) for outcomes $y_{i1}$ and $y_{i2}$. Each point reports the average over 5,000 simulated datasets.
Figure A.5: Power of the difference in CATEs in each covariate ($\lambda_{i}$ and $a_{i}$) for outcomes $y_{i1}$ and $y_{i2}$. Each point reports the average over 5,000 simulated datasets.
Appendix D Generalized Framework with Multiple Mechanisms
@@
$TE^{Y}(z,z^{\prime};X)$ $=Y(z,M_{1}(z;X),...,M_{J}(z;X);X)-Y(z^{\prime},M_{1}(z^{\prime};X),...,M_{J}(z^{\prime};X);X)$ (16)
@@
$DE^{Y}(z,z^{\prime};X)=Y(z,M_{1}(z;X),...,M_{J}(z;X);X)-Y(z^{\prime},M_{1}(z;X),...,M_{J}(z;X);X)$ (17)
$IE^{Y}_{j}(z,z^{\prime};X)=Y$ $(z^{\prime},M_{j-}(z^{\prime};X),M_{j}(z;X),M_{j+}(z;X);X)$- (18)
@@
The direct effect, $DE^{Y}(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediators at potential outcomes $M_{j}(z;X)$. It is not necessary to believe that treatments produce unmediated (direct) effects on outcomes to use this framework. We allow for direct effects in the interest of generality. The indirect effects, measures the effect on the outcome that operates by changing the potential outcome of the mediator. As is standard, we can then re-write the total effect as follows:
$TE^{Y}(z,z^{\prime};X)$ $=DE^{Y}(z,z^{\prime};X)+\sum_{j=1}^{J}IE^{Y}_{j}(z,z^{\prime};X),$ (19)
@@
$ATE^{Y}(z,z^{\prime})$ $=E_{X}[Y(z,M_{j}(z;X);X)-Y(z^{\prime},M_{j}(z^{\prime};X);X)]$ (20)
$=E_{X}[DE^{Y}(z,z^{\prime};X)+\sum_{j=1}^{J}IE^{Y}_{j}(z,z^{\prime};X)]$ (21)
$=ADE^{Y}(z,z^{\prime})+\sum_{j=1}^{J}AIE^{Y}_{j}(z,z^{\prime})$ (22)
Because there are multiple mediators, we generalize Assumption 2 from the main text while maintaining Assumption 1
@@
Given $z,z^{\prime}\in Z$ and $x,x^{\prime}\in X_{k}$, $X_{k}$ is excluded to the indirect effect of any other mechanism, $j^{\prime}\neq j$, $AIE^{Y}_{j^{\prime}}$: $AIE^{Y}_{j^{\prime}}(X_{k}=x)=AIE^{Y}_{j^{\prime}}(X_{k}=x^{\prime})$.
The only difference between Assumption 2 and Assumption 3 is that Assumption 3 is a generalization to Mechanisms $j\in 3,...,J$.